A geometric progression has a second term of 12 and a sum to infinity of 54. Find the possible values of the first term of the progression


1
Expert's answer
2022-03-22T16:37:16-0400

"ar=12, \;\frac{a}{1-r}=54."

"r=\frac{12}{a},\;\frac{a^2}{a-12}=54."

"a^2-54a+648=0."

"a=18,\;a=36."